OpenStudy (bmk614):

A solid has a circular base of radius 3. If every plane cross section perpendicular the the x-axis is an equilateral triangle, then the volume is a) 36 b) 12 sqr3 c) 18 sqr3 d) 24 sqr3 e) 36 sqr3

OpenStudy (dessyj1):

Do you know what your integral is going to look like?

jhonyy9 (jhonyy9):



OpenStudy (bmk614):

Well I got the intergral \[\int\limits_{ }^{ }\pi r^2*\pi r^2 dr\] but I am not too sure about it.

OpenStudy (irishboy123):